OpenStudy (anonymous):

A manufacturer of a product has fixed costs totaling $8295. He finds he can never sell any of the product unless the price is below $97, and for each decrease of $5 in the price, he is able to sell 100 more units. Find the cost function, demand function, revenue function, and profit function

OpenStudy (anonymous):

Each unit costs an additional $47 to produce

OpenStudy (anonymous):

C(x) = 8925 + 47 x , thats all i got, cant do the demand equation

OpenStudy (amistre64):

what do you mean by "each unit costs and additional" ? is that just 45 per unit?

OpenStudy (amistre64):

47 if i read it right lol

OpenStudy (anonymous):

, yes, per item it costs 47 to produce

OpenStudy (anonymous):

, i need help with the demand equation

OpenStudy (anonymous):

fixed costs means like overhead, building and stuff

OpenStudy (anonymous):

rent

OpenStudy (amistre64):

fixed costs = 8295. price < 97 -5 in the price, sell +100 Find the cost function, demand function, revenue function, and profit function Cost(x) = 47x + 8295 Revenue(x) = x*Quantity(x) Quantity(x) = has a slope of 100/-5 = -20 = -20x +20(Px) +Py, if its linear

OpenStudy (amistre64):

Quantity would mean Demand

OpenStudy (amistre64):

how much can they sell at $97 tho

OpenStudy (anonymous):

where is your demand function

OpenStudy (amistre64):

still working on it, but itd be nice to know how much can be sold at a given price

OpenStudy (amistre64):

97,0 maybe

OpenStudy (anonymous):

i guess we start with 97 as the price

OpenStudy (anonymous):

(97,0) and (92, 100)

OpenStudy (amistre64):

Demand(x) = has a slope of 100/-5 = -20 = -20x +20(97) = -20x +194

OpenStudy (anonymous):

let quantity demanded be a function of price

OpenStudy (amistre64):

i dropped a zero :) Demand(x) = -20x+1940

OpenStudy (anonymous):

, wait, where did you get x?

OpenStudy (amistre64):

x is just what im using for price

OpenStudy (anonymous):

i have D = -20 p + 97

OpenStudy (anonymous):

x should stand for units sold

OpenStudy (amistre64):

your wrong on that

OpenStudy (anonymous):

we used x earlier for units sold

OpenStudy (amistre64):

change x to p then :)

OpenStudy (amistre64):

D(p) = -20p +1940

OpenStudy (anonymous):

actually , the units sold are the quantity demanded

OpenStudy (anonymous):

so we have x = -20 p + 97

OpenStudy (amistre64):

not 97

OpenStudy (amistre64):

at p=97 you sell 0

OpenStudy (anonymous):

one sec

OpenStudy (amistre64):

0 = -20(97) + C

OpenStudy (amistre64):

C = 1940

OpenStudy (anonymous):

hmmm, i thought (97,0) is the y intercept

OpenStudy (amistre64):

x intercept :)

OpenStudy (amistre64):

or rather, price intercept

OpenStudy (amistre64):

graphic wise it can be whatever axis you want to put it on

OpenStudy (anonymous):

we have (97,0) and (92,100)

OpenStudy (anonymous):

we are drawing the line through those two points

OpenStudy (amistre64):

slope was given as 100/-5 to begin with = -20

OpenStudy (anonymous):

(price, units sold/quantity demanded)

OpenStudy (amistre64):

the price intercept (97,0)

OpenStudy (anonymous):

so it should be x = -20 p + 97

OpenStudy (anonymous):

oh no

OpenStudy (amistre64):

D(p)= -20p + C , at a price intercept would to the independant variable

OpenStudy (anonymous):

youre right

OpenStudy (amistre64):

thats why i used "x" to show price

OpenStudy (anonymous):

ok so x = -20 p + 1940

OpenStudy (amistre64):

but p is fine, just as long as we know where its going :)

OpenStudy (anonymous):

why did you call it C for price intercept

OpenStudy (amistre64):

Demand(price) = -20(price) + 1940 Revenue(p) = price*Demand(price)

OpenStudy (amistre64):

..C just means some constant that has to be determined

OpenStudy (amistre64):

C is the demand intercept if anything

OpenStudy (anonymous):

ok, thought it was cost

OpenStudy (anonymous):

C is the quantity demanded intercept

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

or 1940

OpenStudy (anonymous):

ok now what do we do

OpenStudy (amistre64):

Now we can determine Revenue(with regards to price)

OpenStudy (amistre64):

R(p) = p*D(p) R(p) = -20p^2 +1940p

OpenStudy (amistre64):

Revenue tells us how much we bring in for a given price

OpenStudy (amistre64):

Profit = Revenue - Costs

OpenStudy (anonymous):

but original problem was in terms of x , not p

OpenStudy (anonymous):

i assume that units sold (x) is equal to quantity demanded

OpenStudy (amistre64):

thats fine

OpenStudy (amistre64):

but x will be a function of p

OpenStudy (anonymous):

otherwise we cant go further

OpenStudy (amistre64):

since quantity depends on price

OpenStudy (anonymous):

right, ill be back in 3 minutes

OpenStudy (anonymous):

can you hold a sec

OpenStudy (anonymous):

unless you have the answer

OpenStudy (amistre64):

Q(p) = -20p +1940 R(p) = -20p^2 +1940p C(Q(p)) = 47(Q(p)) + 8295 = 47(-20p +1940) + 8295 P(p) = -20p^2 +1940p -47(-20p +1940) -8295

OpenStudy (amistre64):

simplify is wanted

OpenStudy (amistre64):

max profits are when p=$72 if its right :)

OpenStudy (anonymous):

hmm, so you did everything in terms of price ?

OpenStudy (anonymous):

interesting, brb

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

since the number of items depends on the number sold; x is a function of p; or rather x = Q(p)

OpenStudy (anonymous):

but is that a fair assumption, the quantity demanded is the same as the number of units sold?

OpenStudy (amistre64):

you dont want to mass produce; that just wastes money if noone will buy

OpenStudy (amistre64):

let the amount that you can sell determine the number you will make

OpenStudy (anonymous):

so the number of units demanded or sold will be the number of units you will produce or supply?

OpenStudy (amistre64):

yes, ideally

OpenStudy (amistre64):

you dont want a shortage or a surplus

OpenStudy (anonymous):

but in wikipedia there is a supply curve and a demand curve

OpenStudy (anonymous):

and they are not the same

OpenStudy (anonymous):

but here we will just assume they are the same ?

OpenStudy (amistre64):

and where the two shall meet is the ideal condition

OpenStudy (amistre64):